OpenStudy (anonymous):

Find the diameter of a cone that has a volume of 29.31 cubic inches and a height of 7 inches. Use 3.14 for pi. 4 inches 2 inches 8 inches 6 inches

Directrix (directrix):

@luka1998 What is the formula for the volume of a right circular cone?

Directrix (directrix):

The volume of a cone can be found by multiplying 1/3 the area of its circular base times the height of the cone.

Directrix (directrix):

V =1/3* pi * r^2 * h where pi*r^2 is the area of the base and h is the height 29.31 = 1/3 * 3.14 * r^2 * 7 @luka1998 Solve the equation for r.

Directrix (directrix):

@luka1998 What is (1/3) * 3.14 * 7 -- > Just multiply those three numbers and post what you get.

Directrix (directrix):

@luka1998

OpenStudy (anonymous):

7.326

Directrix (directrix):

Okay, so 29.31 = 1/3 * 3.14 * r^2 * 7 29.31 = 7.326 * r^2 29.31/7.326 = r^2 So, divide 29.31 by 7.327 and take the square root to get r. Don't forget to double r at the end because the problem asks for diameter.

OpenStudy (anonymous):

4

Directrix (directrix):

The diameter is 4; yes, that is what I got.

OpenStudy (anonymous):

@Directrix THX

Directrix (directrix):

You are welcome.

OpenStudy (compassionate):

\[V = \frac{ 1 }{ 3 }\pi \times r^2h \rightarrow \sqrt{\frac{ 29.31 - 1.05 }{ 7 }} \times 2 = r \rightarrow r = 2, d = r \times 2, d = 4\]